OpenStudy (anonymous):

(-1/2) (3/5) (-2/3) (5/8) , using distributive property

OpenStudy (anonymous):

are they all multiplying each other?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

\[\frac{ -1 }{ 2 } \times \frac{ 3 }{ 5 } \times \frac{ -2 }{ 3 } \times \frac{ 5 }{ 8 }\]

OpenStudy (anonymous):

and then do you have any idea to solve it?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

Not really

OpenStudy (anonymous):

are you still need help?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

 ok?